Alexia has spread her collection of 20 rectangular stamps on a circular table. the table is 4ft in diameter, and each of the stamps is 2 in.x 1 1/2 in. Unfortunatley, a drop of ink from Alexia's pen flies off and lands randomly somewhere on the table. what is the probability that the tink will land on one of her stamps?
Answer by stanbon(75887)

Alexia has spread her collection of 20 rectangular stamps on a circular table. the table is 4ft in diameter, and each of the stamps is 2 in.x 1 1/2 in. Unfortunatley, a drop of ink from Alexia's pen flies off and lands randomly somewhere on the table. what is the probability that the tink will land on one of her stamps?
-----------
Area of one stamp = 2*(3/2)= 3 sq in
Area of 20 stamps = 20*3 sq in = 60 sq in
--------------
Area of circle = (pi)(24 in)^2
------------
Prob (ink lands on stamp) = 60/[24^2 * pi] = 0.033...
Cheers,
Stan H.